### v2

@@ Appendix A Appendix
which is monotone and remains constant on [1,0)[-1,0) and on $(m,m+1]$. By Theorem [36, Theorem 12.12.2], there exists $f_{\infty}^{m}\in D([-1,m+1])$ such that $f_{k}^{m}$ converges to $f_{\infty}^{m}$ in D([1,m+1])D([-1,m+1]) as $k\to\infty$ along some subsequence, which is potentially a further subsequence of the subsequence corresponding to m1m-1. In particular, by [36, Corollary 12.9.1], we can find a continuity point $t_{m}\in(m-1,m]$ of $f_{\infty}^{m}$ such that $f_{k}|_{[-1,t_{m}]}=f_{k}^{m}|_{[-1,t_{m}]}$ converges to $f_{\infty}^{m}|_{[-1,t_{m}]}$ in $D([-1,t_{m}])$ as $k\to\infty$ along that subsequence (note also that $t_{m}\uparrow\infty$ by construction). Combining the diagonal argument, Theorem A.2 and the right-continuity of the limit function, we can find an $f_{\infty}\in D([-1,\infty))$ such that $f_{k}$ converges to $f_{\infty}$ in $D([-1,\infty))$ as $k\to\infty$ along some subsequence. ∎

### v3

@@ Appendix C Remaining Technical Proofs
For each $n\geq 1$, we point out that the unique solution to (3.5) can be obtained by an iteration procedure similar to that in the construction of minimal solutions: $(G,X^{\lambda^{n}},D^{n})=\lim_{N\to\infty}(G,X^{\lambda^{n},N},D^{n,N})$, where
$X_{v}^{\lambda^{n},0}(t)=x_{v}+Z_{v}(t),\quad X_{v}^{\lambda^{n},N+1}(t)=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t,X_{u}^{\lambda^{n},N})$
As
$X_{v}^{\lambda^{n},N+1}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t,X_{u}^{\lambda^{n},N})\geq x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{n,N}\}},$
$0=X_{v}(\tau_{v})=x_{v}+Z_{v}(\tau_{v})\geq x_{v}+Z_{v}(\tau_{v})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}(t,X_{u}^{\lambda})=X_{v}^{\lambda^{n}},$
$x_{v}+Z_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}+\varepsilon_{\Delta}<0.$
by equation (3.6). Therefore,
$X_{v}^{\lambda^{n}}(t_{\Delta})$ $=x_{v}+Z_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t_{\Delta},X_{u}^{\lambda^{n}})$
$\leq x_{v}+Z_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u}^{n},\infty)}(t_{\Delta})+\varepsilon_{\Delta}$
$\leq x_{v}+Z_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}+\varepsilon_{\Delta}<0,$
To get the convergence of the solution paths $X^{\lambda_{n}}$, we note that
$X_{v}^{\lambda^{n}}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(t,X_{u}^{\lambda^{n}})$
$X_{v}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u},\infty)}(t).$
$\leq d_{M_{1}}\Big(\sum_{u\in N_{G}^{-}(v)}c_{uv}\lambda_{uv}^{n}(\cdot,X_{u}^{\lambda^{n}}),\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u}^{n},\infty)}(\cdot)\Big)$
$+d_{M_{1}}\Big(\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u}^{n},\infty)}(\cdot),\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{[\tau_{u},\infty)}(\cdot)\Big)$
$\quad\to 0\quad\text{as}\quad n\to\infty.$
Therefore, we see that $X_{v}^{\lambda_{n}}$ converges almost surely to $X_{v}$ in $\mathcal{D}$ as $n\to\infty$ by combining the above limit with Lemma 2.3 and [34, Lemma 12.7.3]. ∎